# On unknotting fibered positive knots and braids

### v1

@@ 3. Obstructing the unknotting of fibered positive knots along fibered positive knots
The Gordian distance of two knots is the minimum number of crossing changes required to transform one knot into the other. In the proof of Theorem 1.3 we make use of a new lower bound for Gordian distances in terms of the homology of branched coverings, which we give in Proposition 3.2. For a knot $K$ and an integer $n\geq 2$, denote by $\Sigma_{n}(K)$ the $n$-fold branched covering of $S^{3}$ along $K$ (see e.g. [BZH14, Section 8.E] for an introduction). The action of the deck transformation group $\langle t\,|\,t^{n}\rangle$ on $\Sigma_{n}(K)$ makes the homology groups of $\Sigma_{n}(K)$ into modules over the ring $\mathbb{Z}[t]/(t^{n}-1)$. Since $(1+\dots+t^{n-1})$ acts trivially on homology, one may also consider homology as $R_{n}$-modules, for $R_{n}\coloneqq\mathbb{Z}[t]/(1+\dots+t^{n-1})$. To compute $H_{1}(\Sigma_{n}(K);R_{n})$, we will use the following lemma.
@@
Let the integer $p\geq 2$ be a prime such that $R_{p}$ is a PID.  [^1] This is the case if and only if $p\leq 19$ (see e.g. [Was97, Theorem 11.1]). We expect that an adapted version of Proposition 3.2 holds for all primes $p$, using that $R_{p}$ is always a Dedekind domain. But since we will only use the case $p=3$ in this paper, we do not pursue this further. Let $q\in R_{p}$ be a prime element and $e\geq 1$ an integer. For a knot $K$, denote by $n_{p,q,e}(K)\geq 0$ the maximal integer such that $H_{1}(\Sigma_{p}(K);R_{p})$ admits a summand $R_{p}/q^{f_{1}}\oplus\dots\oplus R_{p}/q^{f_{n_{p,q,e}(K)}}$ with $f_{1},\dots,f_{n_{p,q,e}(K)}\geq e$. Then $|n_{p,q,e}(K)-n_{p,q,e}(J)|$ is a lower bound for the Gordian distance between the knots $K$ and $J$.
Let us first deduce Theorem 1.3 from Proposition 3.2, and prove the proposition afterwards.
@@
$H_{1}(\Sigma_{3}(K11n183);R_{3})$ $\ \cong\ (R_{3}/4)^{\oplus 2},$
$H_{1}(\Sigma_{3}(T(2,3)^{\#2});R_{3})$ $\ \cong\ (R_{3}/2)^{\oplus 2},$
@@
for some integer column vectors $v,w$ and for $x\in\mathbb{Z}$ [Sae99]. Consider the matrix $E$, which has the same size as matrix $C$ but with all entries being zero except for the lower right corner, which is $\mp 1$. Now, let $\widetilde{B}=C+E$. This matrix $\widetilde{B}$ is S-equivalent to the matrix $B$. Define $B^{\prime}=t\widetilde{B}-\widetilde{B}^{\top}$, $C^{\prime}=tC-C^{\top}$, and $E^{\prime}=tE-E^{\top}$, all considered as matrices over the ring $R_{p}$. Note that by Lemma 3.1, $B^{\prime}$ and $C^{\prime}$ serve as presentation matrices for $H_{1}(\Sigma_{p}(K);R_{p})$ and $H_{1}(\Sigma_{p}(J);R_{p})$, respectively. Furthermore, we have the relationship $C^{\prime}=B^{\prime}-E^{\prime}$, and the rank $\operatorname{rk}_{R_{p}}E^{\prime}$ of the matrix $E^{\prime}$ is equal to $1$. Here, the rank of a matrix (with entries in any commutative ring) is defined as the smallest number of generators for the submodule generated by the column vectors of the matrix.
The rest of the proof applies universally to any PIDs and does not depend on our specific geometric situation. Since $n_{p,q,e}(K)$ and $n_{p,q,e}(J)$ are determined by matrices $B^{\prime}$ and $C^{\prime}$, we refer to them by $n_{p,q,e}(B^{\prime})$ and $n_{p,q,e}(C^{\prime})$, respectively. Our first step now is to prove that
$n_{p,q,e}(C^{\prime})=m-\operatorname{rk}_{S}C^{\prime},$
where $m$ is the size of the matrix $C^{\prime}$, and $S$ is the ring defined as $R_{p}/(q^{e})$. We can simplify the situation by applying the Smith normal form. Because $R_{p}$ is a PID, we can find invertible matrices $U$ and $V$ such that the product $UC^{\prime}V$ results in a diagonal matrix with entries $d_{1},\ldots,d_{n}$, where each $d_{i}$ divides $d_{i+1}$. It is important to note that the rank and the isomorphism type of the cokernel remain unchanged when we multiply by $U$ and $V$. So, $n_{p,q,e}(C^{\prime})$ is the number of $d_{i}$ divisible by $q^{e}$. When we consider $UC^{\prime}V$ as a matrix over the ring $S=R_{p}/(q^{e})$, these entries $d_{i}$ become 00. This proves $\operatorname{rk}_{S}C^{\prime}\leq m-n_{p,q,e}(C^{\prime})$.
Let us now show the reverse inequality. Let $\ell=m-n_{p,q,e}(C^{\prime})$ and for $i\in\{1,\ldots,\ell\}$, let $f_{i}$ be maximal such that $q^{f_{i}}$ divides $d_{i}$. By assumption, we have $0\leq f_{i}<e$. Note that the $S$-module generated by the columns of $UC^{\prime}V$ is isomorphic to $S/q^{e-f_{1}}\oplus\dots\oplus S/q^{e-f_{\ell}}$. Let a system of $a$ generators of this module be given, i.e. a surjective $S$-homomorphism
$S^{a}\to S/q^{e-f_{1}}\oplus\dots\oplus S/q^{e-f_{\ell}}.$
Composing with multiplication by $q^{e-f_{i}-1}$ in the $i$-th component, this yields a surjective homomorphism $S^{a}\to(S/q)^{\oplus\ell}$. Since the target module is annihilated by $q$, this induces a surjective $S/(q)$-homomorphism $(S/(q))^{a}\to(S/q)^{\oplus\ell}$. But since $S/(q)\cong R_{p}/(q)$ is a field, we have $a\geq\ell$. Thus we have established that any system of generators of the column space of $UC^{\prime}V$ over $S$ has at least $\ell$ generators, and so the desired inequality $\operatorname{rk}_{S}C^{\prime}\geq\ell=m-n_{p,q,e}(C^{\prime})$ follows.
To conclude the proof of the proposition, note that it follows quickly from its definition that $\operatorname{rk}_{S}$ is subadditive, and therefore
$n_{p,q,e}(C^{\prime})=m-\operatorname{rk}_{S}C^{\prime}\geq m-\operatorname{rk}_{S}B^{\prime}-\operatorname{rk}_{S}E^{\prime}=n_{p,q,e}(B^{\prime})-1,$
implying $n_{p,q,e}(B^{\prime})-n_{p,q,e}(C^{\prime})\leq 1$. Switching the roles of $B^{\prime}$ and $C^{\prime}$ gives us $|n_{p,q,e}(B^{\prime})-n_{p,q,e}(C^{\prime})|\leq 1$, as desired. ∎
@@ 5. Fibered positive knots with potentially large unknotting number
To see that the minimal number of crossing changes needed to transform $D$ into a diagram of the unknot is $9$, we have checked all possible combinations of at most $8$ crossing changes in $D$ and observe that the Alexander polynomials of the resulting knots are non-trivial. The computations can be accessed at [${}^{+}$]. An unknotting sequence with 9 crossings is easily found in $D$ (see also Proposition 5.3(4) below). ∎
@@
Since we cannot find 8 crossing changes that unknot $\mathcal{K}$, we conjectured (see 1.6) that in fact $u(\mathcal{K})=9$, which would make $\mathcal{K}$ a counterexample to Stoimenow’s 1.1. Let us now discuss why $u(\mathcal{K})>7$ is hard to show. Although the literature is awash with lower bounds for the unknotting number, most of them fall into one of the following three types:
@@
(2) To see that $\mathcal{K}_{1}$ is not braid positive, we compute in [${}^{+}$] Ito’s normalized version of the HOMFLYPT polynomial $\tilde{P}_{\mathcal{K}_{1}}(\alpha,z)$ and observe that it has negative coefficients, failing a condition for braid positive knots [Ito22, Theorem 1.1]. In particular, we observe that $\tilde{P}_{\mathcal{K}_{1}}(\alpha,0)=-\alpha^{4}+7\alpha^{3}+9\alpha^{2}$. Applying the skein relation for the HOMFLYPT polynomial inductively to the $3n$ and $4n$ crossings involved in the crossing changes in Figure 10(b) and (c), we can inductively reduce the general case to the case $n=1$. In particular, it follows that
@@
which has negative coefficients and thus $\mathcal{K}_{n}$ is not braid positive.
@@
The lower bounds in (5) and (6) are given by the Montesinos trick: the double branched cover $\Sigma_{2}(K)$ of $K$ can be obtained from $S^{3}$ by surgery on a link with $u_{q}(K)$ components [MZ23]. Moreover, $\Sigma_{2}(K)$ can be obtained from an integer homology sphere by surgery on a link with $u_{a}(K)$ components (since a knot with Alexander polynomial 1 has a $\mathbb{Z}HS^{3}$ as double branched covering). In particular, the minimal numbers of required generators of $H_{1}(\Sigma_{2}(K);\mathbb{Z})$ is a lower bound for $u_{a}(K)$ and $u_{q}(K)$. To compute the first homology of the double branched cover we recall that the diagram $D_{n}$ consists of the cyclic composition of $12n+$ 1 copies of the tangle $T$. From that, it is not hard to see that the double branched cover of $\mathcal{K}_{n}$ is obtained by taking $S^{1}\times(S^{2}\setminus\bigcup_{2n+1}\mathring{D}^{2})$ and gluing $12n+$ 1 copies of the double branched cover of $T$ to it. On the other hand, the double branched cover of $T$ can be seen to be the figure eight knot sister $m003$, a $1$-cusped hyperbolic manifold realizing the minimum of volumes among such manifolds. (The other minimizer is the figure eight knot complement.) $m003$ can be seen as the complement of a nullhomologous knot in the lens space $L(5,1)$ and thus has homology $\mathbb{Z}\oplus\mathbb{Z}_{5}$. A surgery description of $m003$ is given by performing a $(-5)$-surgery on one of the components of the Whitehead link. (From that surgery description all the above statements are easy to verify.) Now a simple application of the Mayer–Vietoris sequence yields

### v2

@@ 3. Obstructing the unknotting of fibered positive knots along fibered positive knots
The Gordian distance of two knots is the minimum number of crossing changes required to transform one knot into the other. In the proof of Theorem 1.3 we make use of a new lower bound for Gordian distances in terms of the homology of branched coverings, which we give in Proposition 3.5. As preparation, let us make two definitions and prove one lemma in a rather general algebraic context.
Definition 3.1
Let $R$ be a PID, $q\in R$ a prime element and $e\geq 1$ an integer. For a matrix $A$ over $R$, denote by $n_{q,e}(A)$ the maximal integer such that the cokernel of $A$ admits a summand $R/(q^{f_{1}})\oplus\dots\oplus R/(q^{f_{n_{q,e}(A)}})$ with $f_{1},\dots,f_{n_{q,e}(A)}\geq e$.
Definition 3.2
The rank $\operatorname{rk}_{S}A$ of an $m\times m$ matrix $A$ over a commutative ring $S$ is the smallest number of generators for the submodule of $S^{m}$ generated by the columns of $A$.
Lemma 3.3
Let $A$ be an $m\times m$ matrix over a PID $R$, $q\in R$ a prime element and $e\geq 1$ an integer. Let $S$ be the ring $R/(q^{e})$. Then $n_{q,e}(A)=m-\operatorname{rk}_{S}A$.
Proof.
We can simplify the situation by applying the Smith normal form. Because $R$ is a PID, there are invertible matrices $U$ and $V$ such that the product $UAV$ is a diagonal matrix with entries $d_{1},\ldots,d_{m}$, where each $d_{i}$ divides $d_{i+1}$. It is important to note that the rank and the isomorphism type of the cokernel of $A$ remain unchanged when we multiply by $U$ and $V$, i.e. $\operatorname{rk}_{S}A=\operatorname{rk}_{S}UAV$ and $n_{q,e}(A)=n_{q,e}(UAV)$. Hence $n_{q,e}(A)$ equals the number of $d_{i}$ divisible by $q^{e}$. When we consider $UAV$ as a matrix over the ring $S$, these entries $d_{i}$ become 00. This proves $\operatorname{rk}_{S}A\leq m-n_{q,e}(A)$.
Let us now show the reverse inequality. Let $\ell=m-n_{q,e}(A)$ and for $i\in\{1,\ldots,\ell\}$, let $f_{i}$ be maximal such that $q^{f_{i}}$ divides $d_{i}$. By assumption, we have ${0\leq f_{i}<e}$. Note that the $S$-module generated by the columns of $UAV$ is isomorphic to $S/(q^{e-f_{1}})\oplus\dots\oplus S/(q^{e-f_{\ell}})$. Let a system of $a$ generators of this module be given, i.e. a surjective $S$-homomorphism
$S^{a}\to S/(q^{e-f_{1}})\oplus\dots\oplus S/(q^{e-f_{\ell}}).$
Composing with multiplication by $q^{e-f_{i}-1}$ in the $i$-th component, this yields a surjective homomorphism $S^{a}\to(S/(q))^{\oplus\ell}$. Since the target module is annihilated by $q$, this induces a surjective $S/(q)$-homomorphism $(S/(q))^{a}\to(S/(q))^{\oplus\ell}$. But since $S/(q)\cong R_{p}/(q)$ is a field, we have $a\geq\ell$. Thus we have established that any system of generators of the column space of $UAV$ over $S$ has at least $\ell$ generators, and so the desired inequality $\operatorname{rk}_{S}A\geq\ell=m-n_{q,e}(A)$ follows. ∎
We shall now apply our newly found algebraic wisdom to the homology of branched coverings. Namely, for a knot $K$ and an integer $n\geq 2$, denote by $\Sigma_{n}(K)$ the $n$-fold branched covering of $S^{3}$ along $K$ (see e.g. [BZH14, Section 8.E] for an introduction). The action of the deck transformation group $\langle t\,|\,t^{n}\rangle$ on $\Sigma_{n}(K)$ makes the homology groups of $\Sigma_{n}(K)$ into modules over the ring $\mathbb{Z}[t]/(t^{n}-1)$. Since $(1+\dots+t^{n-1})$ acts trivially on homology, one may also consider homology as $R_{n}$-modules, for $R_{n}\coloneqq\mathbb{Z}[t]/(1+\dots+t^{n-1})$. To compute $H_{1}(\Sigma_{n}(K);R_{n})$, we will use the following lemma.
@@
Let the integer $p\geq 2$ be a prime such that $R_{p}$ is a PID.  [^1] This is the case if and only if $p\leq 19$ (see e.g. [Was97, Theorem 11.1]). We expect that an adapted version of Proposition 3.5 holds for all primes $p$, using that $R_{p}$ is always a Dedekind domain. But since we will only use the case $p=3$ in this paper, we do not pursue this further. Let $q\in R_{p}$ be a prime element and $e\geq 1$ an integer. For a knot $K$, denote by $n_{p,q,e}(K)\geq 0$ the maximal integer such that $H_{1}(\Sigma_{p}(K);R_{p})$ admits a summand $R_{p}/(q^{f_{1}})\oplus\dots\oplus R_{p}/(q^{f_{n_{p,q,e}(K)}})$ with $f_{1},\dots,f_{n_{p,q,e}(K)}\geq e$. Then $|n_{p,q,e}(K)-n_{p,q,e}(J)|$ is a lower bound for the Gordian distance between the knots $K$ and $J$.
By Lemma 3.4, $n_{p,q,e}(K)$ equals $n_{q,e}(tB-B^{\top})$ (see Definition 3.1) for a Seifert matrix $B$ of $K$. Let us first deduce Theorem 1.3 from Proposition 3.5, and prove the proposition afterwards.
@@
$H_{1}(\Sigma_{3}(K11n183);R_{3})$ $\ \cong\ (R_{3}/(4))^{\oplus 2},$
$H_{1}(\Sigma_{3}(T(2,3)^{\#2});R_{3})$ $\ \cong\ (R_{3}/(2))^{\oplus 2},$
@@
for some integer column vectors $v,w$ and for $x\in\mathbb{Z}$ [Sae99]. Consider the matrix $E$, which has the same size as matrix $C$ but with all entries being zero except for the lower right corner, which is $\mp 1$. Now, let $\widetilde{B}=C+E$. This matrix $\widetilde{B}$ is S-equivalent to the matrix $B$. Define $B^{\prime}=t\widetilde{B}-\widetilde{B}^{\top}$, $C^{\prime}=tC-C^{\top}$, and $E^{\prime}=tE-E^{\top}$, all considered as matrices over the ring $R_{p}$. Note that by Lemma 3.4, $B^{\prime}$ and $C^{\prime}$ serve as presentation matrices for $H_{1}(\Sigma_{p}(K);R_{p})$ and $H_{1}(\Sigma_{p}(J);R_{p})$, respectively. Furthermore, we have the relationship $C^{\prime}=B^{\prime}-E^{\prime}$. With $S=R_{p}/(q^{e})$, the rank $\operatorname{rk}_{S}E^{\prime}$ of the matrix $E^{\prime}$ is equal to $1$ (see Definition 3.2). We now have
$n_{p,q,e}(J)=m-\operatorname{rk}_{S}C^{\prime}\geq m-\operatorname{rk}_{S}B^{\prime}-\operatorname{rk}_{S}E^{\prime}=n_{p,q,e}(K)-1,$
where the equalities are due to Lemma 3.3, and the inequality due to the subadditivity of $\operatorname{rk}_{S}$, which follows quickly from its definition. Hence $n_{p,q,e}(K)-n_{p,q,e}(J)\leq 1$. Switching the roles of $K$ and $J$ gives us $|n_{p,q,e}(J)-n_{p,q,e}(J)|\leq 1$, as desired. ∎
@@ 5. Fibered positive knots with potentially large unknotting number
To see that the minimal number of crossing changes needed to transform $D$ into a diagram of the unknot is $9$, we have checked all possible combinations of at most $8$ crossing changes in $D$ and observed that the Alexander polynomials of the resulting knots are non-trivial. The computations can be accessed at [KLM+]. An unknotting sequence with 9 crossings is easily found in $D$ (see also Proposition 5.3(4) below). ∎
@@
Since we cannot find 8 crossing changes that unknot $\mathcal{K}$, we conjectured (see 1.6) that in fact $u(\mathcal{K})=9$, which would make $\mathcal{K}$ a counterexample to Stoimenow’s 1.1. Let us now discuss why $u(\mathcal{K})>7$ is hard to show. Although the literature is awash with lower bounds for the unknotting number, most of them fall into one of the following four types:
@@
(2) To see that $\mathcal{K}_{1}$ is not braid positive, we compute in [KLM+] Ito’s normalized version of the HOMFLYPT polynomial $\tilde{P}_{\mathcal{K}_{1}}(\alpha,z)$ and observe that some of its coefficients are negative, failing a condition for braid positive knots [Ito22, Theorem 1.1]. In particular, we observe that $\tilde{P}_{\mathcal{K}_{1}}(\alpha,0)=-\alpha^{4}+7\alpha^{3}+9\alpha^{2}$. Applying the skein relation for the HOMFLYPT polynomial inductively to the $3n$ and $4n$ crossings involved in the crossing changes in Figure 10(b) and (c), we can inductively reduce the general case to the case $n=1$. In particular, it follows that
@@
which has a negative coefficient and thus $\mathcal{K}_{n}$ is not braid positive.
@@
The lower bounds in (5) and (6) are given by the Montesinos trick: the double branched cover $\Sigma_{2}(K)$ of $K$ can be obtained from $S^{3}$ by surgery on a link with $u_{q}(K)$ components [MZ23]. Moreover, $\Sigma_{2}(K)$ can be obtained from an integer homology sphere by surgery on a link with $u_{a}(K)$ components (since a knot with Alexander polynomial 1 has a $\mathbb{Z}HS^{3}$ as double branched covering). In particular, the minimal numbers of required generators of $H_{1}(\Sigma_{2}(K);\mathbb{Z})$ is a lower bound for $u_{a}(K)$ and $u_{q}(K)$. To compute the first homology of the double branched cover we recall that the diagram $D_{n}$ consists of the cyclic composition of $12n+$ 1 copies of the tangle $T$. From that, it is not hard to see that the double branched cover of $\mathcal{K}_{n}$ is obtained by taking $S^{1}\times(S^{2}\setminus\bigcup_{2n+1}\mathring{D}^{2})$ and gluing $12n+$ 1 copies of the double branched cover of $T$ to it. On the other hand, the double branched cover of $T$ can be seen to be the figure eight knot sister $m003$, a $1$-cusped hyperbolic manifold realizing the minimum of volumes among such manifolds. (The other minimizer is the figure eight knot complement.) $m003$ can be seen as the complement of a nullhomologous knot in the lens space $L(5,1)$ and thus has homology $\mathbb{Z}\oplus\mathbb{Z}_{5}$. In SnapPy [CDGW], Manifold(’L5a1(-5,1)(0,0)’).identify() shows that the $(-5)$-surgery on one of the components of the Whitehead link $L5a1$ is diffeomorphic to $m003$. (From that surgery description all the above statements are easy to verify.) Now a simple application of the Mayer–Vietoris sequence yields